OpenStudy (omarbirjas):

Sketch the graph of each function in the interval from 0 to 2pi

OpenStudy (omarbirjas):

\[y=2\cos(\frac{ \pi }{ 2})\theta\]

OpenStudy (campbell_st):

what do you know about the curve? amplitude and period..?

OpenStudy (omarbirjas):

how many cycles it makes from 0 to 2pi

OpenStudy (campbell_st):

ok...a standard form is \[y = a \cos (bx) \] a = amplitude... so in your question a = 2 the value of b in your question is \[b = \frac{\pi}{2}\] and the period is found by \[period = \frac{2\pi}{b}\] to to find the period in your curve its \[period = \frac{\pi}{\frac{\pi}{2}}\] so can you simplify it..?

OpenStudy (omarbirjas):

1/2

OpenStudy (campbell_st):

oops... the period is \[period = \frac{2\pi}{\frac{\pi}{2}}\] sorry for the mistake

OpenStudy (omarbirjas):

1

OpenStudy (omarbirjas):

no probs

OpenStudy (campbell_st):

well you are dividing by a fraction so flip and multiply \[2\pi \times \frac{2}{\pi} = 4\] so over a distance of 4, the curve completes 1 cycle

OpenStudy (omarbirjas):

what? how does it equal 4?

OpenStudy (campbell_st):

the period \[period = \frac{2\pi}{\frac{\pi}{2}} \] I think you'll find that the period is 4 if you simplify it.

OpenStudy (omarbirjas):

Well i put it in a calc and got 1

OpenStudy (campbell_st):

if you would like to check you can use https://www.desmos.com/calculator it's an excellent online graphing site

OpenStudy (omarbirjas):

i tried it....it gave me a straight line...

OpenStudy (campbell_st):

replace theta with x... and you will get the curve

OpenStudy (campbell_st):

and if you feel the period is 1... then that's what you should use.

OpenStudy (omarbirjas):

i just want to know how its 4

OpenStudy (campbell_st):

\[\frac{2\pi}{\frac{\pi}{2}} = 2\pi \div(\pi \div 2) = 4 \] the rule for dividing by a fraction 1. find the reciprocal of the fraction. 2. multiply

OpenStudy (campbell_st):

so 1. reciprocal is \[\frac{2}{\pi}\] 2. multiply \[2\pi \times \frac{2}{\pi} = 4\]

OpenStudy (omarbirjas):

why do we reciprocate?

OpenStudy (campbell_st):

that's the rule for dividing by fractions... the question is asking \[how~~many~~~\frac{\pi}{2}~~~ make ~~2\pi\] and you need 4

OpenStudy (omarbirjas):

ohhh ok

OpenStudy (omarbirjas):

gotcha...proceed please

OpenStudy (campbell_st):

so now you know the period and amplitude you can draw the graph.  its a rough sketch... but I hope it makes sense